A particle P is projected from a point I with an initial velocity of 70m/s at an angle of 30° to the horizontal. At the same instant, a second particle Q is projected in the same direction, 20m from point I with initial speed 25m/s from a point level with O. If particle Q is displaced to 90m from point of release C. If the particle Q collides with P at 80m from the point of release C, find the angle of projection of Q and find when the collision occurs.



1
Expert's answer
2021-11-14T17:13:29-0500

for projectile:

"x=x_0+v_{0x}t"

"y=y_0+v_{0y}t-gt^2\/2"


for particle P:

"x_P=70tcos30\degree"

"y_P=70tsin30\degree-4.9t^2"


for particle Q:

"x_Q=25tcos\alpha+20"

"y_Q=25tsin\alpha-4.9t^2"


If the particles Q collide with P, then:

"y_P=y_Q"

"70sin30\degree=25sin\alpha"

"sin\alpha=70sin30\degree/25=7/5>1"


then:

"y_Q=25tsin\alpha+4.9t^2" (Q is projected in -y direction)


for collision:

"70tsin30\degree-4.9t^2=25tsin\alpha+4.9t^2"

"70sin30\degree-9.8t=25sin\alpha"


"x_Q^2+y_Q^2=(25tsin\alpha+4.9t^2)^2+(25tcos\alpha+20)^2=80^2"


"625t^2+9.8t^3(35-9.8t)+24t^4+1000tcos\alpha+400=6400"


"x_P=x_Q"

"25tcos\alpha+20=70tcos30\degree"


then:


"625t^2+9.8t^3(35-9.8t)+24t^4+40(35t\sqrt 3-20)=6000"

"-72t^4+343t^3+625t^2+2425t-6800=0"


time of collision:

"t_1=1.66" s, "t_2=6.54" s

"sin\alpha_1=(35-9.8\cdot1.66)/25=0.749"

"\alpha_1=48.5\degree"

"sin\alpha_1=(35-9.8\cdot6.54)/25=-1.16<-1"


so, we have:

angle of projection of Q:

"\alpha_1=-48.5\degree" (since Q is projected in -y direction)

time of collision:

"t=1.66" s